# ContourFormer:Real-Time Contour-Based End-to-End Instance Segmentation Transformer

### v2

@@ Abstract
11footnotetext: Email:yaoww1@csrzic.com.22footnotetext: code will be made publicly available at: https://github.com/talebolano/Contourformer
@@ 4.2 SBD
Contourformer - 67.9 75.9 64.0
$\mathrm{Contourformer}^{+}$ 69.2 77.2 65.7
@@ 4.3 COCO
Contourformer - HGNetv2-B2 36.1 36.5
$\mathrm{Contourformer}^{+}$ HGNetv2-B3 37.2 37.6
@@ 4.5 Ablation Study
Building on this foundation of $N_{c}=8$, we further investigated two additional aspects: whether employing sub-contour bounding boxes as deformable attention sampling ranges would influence model accuracy, and the effect of contour fine-grained distribution refinement on model precision. As demonstrated in Table V, using sub-contour bounding boxes as deformable attention sampling ranges improved $\mathrm{AP}_{vol}$ by 1.3, without adversely affecting inference time or memory usage. Furthermore, incorporating contour fine-grained distribution refinement yielded an additional improvement of 0.8 in $\mathrm{AP}_{vol}$, albeit with a notable increase in inference time.These experiments provided valuable insights into the trade-offs between different components and parameter choices, guiding us toward an optimal configuration that maximizes performance while considering computational efficiency.

### v3

@@ Abstract
11footnotetext: Email:yaoww1@csrzic.com.22footnotetext: The code is available at: https://github.com/talebolano/Contourformer
@@ 4.2 SBD
Contourformer - 67.6 75.5 63.6
$\mathrm{Contourformer}^{+}$ 69.2 76.9 65.8
@@ 4.3 COCO
Contourformer - HGNetv2-B2 36.4 36.5
$\mathrm{Contourformer}^{+}$ HGNetv2-B3 37.3 37.6
@@ 4.5 Ablation Study
Building on this foundation of $N_{c}=8$, we further investigated two additional aspects: whether employing sub-contour bounding boxes as deformable attention sampling ranges would influence model accuracy, and the effect of contour fine-grained distribution refinement on model precision. As demonstrated in Table V, using sub-contour bounding boxes as deformable attention sampling ranges improved $\mathrm{AP}_{vol}$ by 1.3, without adversely affecting inference time or memory usage. Furthermore, incorporating contour fine-grained distribution refinement yielded an additional improvement of 0.5 in $\mathrm{AP}_{vol}$, albeit with a notable increase in inference time.These experiments provided valuable insights into the trade-offs between different components and parameter choices, guiding us toward an optimal configuration that maximizes performance while considering computational efficiency.